OpenStudy (anonymous):

Evaluate the expression 6y -x^{2} for x = 2, y = 5. A. 784 B. 26 C. 324 D. 54

OpenStudy (anonymous):

@LegendarySadist

OpenStudy (anonymous):

\[6y -x^{2} for x = 2, y = 5.\]

OpenStudy (anonymous):

So \[\large \sf 6(5)-2^{2}\]

OpenStudy (anonymous):

2x2? or 2x2x2?

OpenStudy (anonymous):

The exponent is the number of times you multiply the base by itself. Since the exponent is 2, you multiply the base by itself twice. So 2x2.

OpenStudy (anonymous):

6x5=30 2x2 = 4 30 - 4 = 26 So 26 is the answre

OpenStudy (anonymous):

That would be correct

OpenStudy (anonymous):

Hi @mathmale

OpenStudy (anonymous):

owo?

OpenStudy (mathmale):

Yes, it's correct. 6x5=30 2x2 = 4 30 - 4 = 26 has resulted in the correct answer, but I have a problem with your having more than 1 equal sign in one equation.

OpenStudy (anonymous):

What?

OpenStudy (mathmale):

Just do the math from top down, not from left to right. 6*5 = 30. Yes. 2*2 = 4 Yes. 30 - 4 = 26.

OpenStudy (mathmale):

6x5=30 2x2 = 4 30 - 4 = 26 would be confusing to many readers.

OpenStudy (anonymous):

o

OpenStudy (anonymous):

Didn't know that

OpenStudy (anonymous):

owo. Thank you

OpenStudy (mathmale):

Surely hope you'll take this suggestion in the spirit of helpfulness that was intended. You're welcome! Take care.